# The non-Abelian geometry, topology, and dynamics of a nonreciprocal Su-Schrieffer-Heeger ladder

### v1

@@ IV Non-Abelian dynamics under an external constant force in the pseudo-Hermitian symmetric regime
$\equiv$ $W_{k_{0}\rightarrow k}\alpha(0)=U_{k}^{-1}U_{k_{0}},$
where
[…]
is a four-band Wilson line for non-Hermitian systems along the path from $k_{0}$ to $k$ in $k$-space [50]. Equation (28) means that the evolution only depends on the start and the final points, regardless of the path’s selection, as shown in Fig. 2(h), which reproduces the evolution by Eq. (26) under a very strong force, as shown in Fig. 2(g).
When $w_{2}\ll F\ll\Delta$, where $w_{2}$ is the maximum energy difference between the lowest/highest two bands and $\Delta$ is the minimum energy difference (i.e., the gap) between the middle two bands, the lowest/highest two bands can be approximated as being degenerate in the whole BZ and the excitation to the other two bands can be ignored, that is, the dynamics can be considered as an adiabatic evolution in the subspace of the lowest/highest two bands. Thus, Eq. (26) can be approximately decoupled into two independent evolutions in each subspace:
[…]
$\equiv$ $W^{(\pm)}_{k_{0}\rightarrow k}\alpha^{(\pm)}(0),$
where (±)plus-or-minus($\pm$)( ± ) means the upper/lower $2\times 2$ diagonal blocks in the labeled matrices and $2\times 1$ blocks in labeled column vectors. In this case, the non-Abelian dynamics is well described by the two-band Wilson lines for non-Hermitian systems [50] in corresponding subspaces:
[…]
as shown in Fig. 2(i), which is similar to the evolution by Eq. (26) under a medium force, as shown in Fig. 2(c).
The non-Abelian dynamics for other strengths of the force cannot be well captured by Wilson lines, as shown in other cases of Fig. 2. The details of the numerical method for calculating the two Wilson lines can be found in Appendix D.
V Bulk-boundary correspondence and the critical NHSE
Figure 3: (a) Comparison between the winding number $w$ (black solid line) under PBCs and the open-bulk winding number $w_{r}$ (red crosses) under OBCs along the orange dotted line at $\chi=2$ in Fig. 1(b), where the parameters for calculating $w_{r}$ are $(L,L^{\prime},l)=(200,160,20)$. (b) The spectrum flow in complex energy plane under OBCs for various unit-cell numbers from $L=5$ to $100$ (from dark blue to light green dots) at weak inter-leg coupling $\chi=0.01$ of the positive case. The spectrum (two red solid lines with dots in between) under OBCs without inter-leg coupling for $L=100$ and the spectrum (gray dashed curve) under PBCs at $\chi=0.01$ of the positive case are plotted for reference. Insets: Typical density distributions of bulk and end eigenstates in leg-$\uparrow$ (purple) and leg-$\downarrow$ (yellow) for several system sizes indicated by the arrows. The other parameters are $(\kappa,\delta)=(2.5,2)$. (c) The averaged density distributions versus the scaled unit-cell index for $L=5$ (circles), $50$ (triangles), and $100$ (squares) in leg-$\uparrow$ (purple) and leg-$\downarrow$ (yellow).
To investigate whether the ladder satisfies the BBC, we numerically calculate the spectra under OBCs with $L=400$ unit cells, as shown in top layers of Figs. 1(c)-1(r), and find that except for the topological zero-mode end states, which are protected by the chiral symmetry, the bulk spectra are identical to those under PBCs (the bottom layers in Figs. 1(c)-1(r)), indicating the survival of the BBC in this ladder. In addition, we also numerically calculate the open-bulk winding number in real space under OBCs [58, 50]:
where $\hat{C}$ is the chiral operator, $\hat{Q}\equiv\sum_{n}\big{[}|\phi_{n}^{(r)}\rangle\langle\phi_{n}^{(l)}|-\hat{C}|\phi_{n}^{(r)}\rangle\langle\phi_{n}^{(l)}|\hat{C}\big{]}$ is the “$Q$-matrix” operator constructed by the $n$th chiral-operation paired right/left eigenstates $\big{\{}|\phi_{n}^{(r,l)}\rangle,\hat{C}|\phi_{n}^{(r,l)}\rangle\big{\}}$ of $\hat{H}_{\text{NH}}$ under OBCs, and the primed trace means that the trace is done over a central $L^{\prime}=L-2l$ unit cells out of the whole $L$ unit cells of the ladder with $l$ unit cells being cut off from each end of the ladder when the numerical calculation is executed to avoid the boundary effects. The numerical results in Fig. 3(a) show that $w_{r}$ is close to $w$ for large systems, which hints that there is no BBC breaking and thus no skin effect in the thermodynamic limit $L\rightarrow\infty$. This can also be reflected by the winding number for energy under PBCs [22, 59]:
However, when the inter-leg coupling vanishes (i.e., $\chi=0$), the ladder is decoupled into two independent nonreciprocal SSH chains, which are well known for the breakdown of BBC and the emergence of NHSEs under OBCs. This indicates that in the thermodynamic limit, the BBC is abruptly restored with the introduction of an infinitesimal inter-leg coupling, which is called the critical NHSE [42]. The critical NHSE leads to a noticeable alteration of the energy spectra and eigenstates under OBCs in the weak inter-leg coupling limit as the system size increases, as shown in Fig. 3(b). When the size is small, the energy spectra and eigenstates are close to those of the decoupled nonreciprocal SSH chains under OBCs, i.e., the spectra are curve/line-like and the eigenstates demonstrate NHSEs, indicating that the coupling effects are negligible; when the size becomes larger, they converge to those under PBCs, i.e., the curve/line-like spectra expand to loops and the eigenstates tend to be delocalized without NHSEs as Bloch waves, indicating that the weak coupling becomes stronger in the thermodynamic limit. Quantitatively, for a finite-size weakly inter-leg coupled ladder under OBCs, the distribution of bulk eigenstates remain the exponential form $|\phi_{j}|^{2}\propto e^{-|j-j_{b}|/\xi}$ of the skin bulk states of the nonreciprocal SSH chain, localized in the boundary unit cell $j_{b}$, but the decaying length $\xi$ of the amplitude $\phi_{j}$ is size-dependent and satisfies the scale-free power law [42, 43]: $\xi(L)\propto L^{c}$, unlike the nonreciprocal SSH chain, where $\xi^{-1}=\frac{1}{2}\ln|\kappa_{+}/\kappa_{-}|$ is size-independent [18]. To show the power law, we define the averaged density distributions in leg-$\sigma$ ($\sigma=\uparrow$ and $\downarrow$) for certain system size with unit-cell number $L$:
where the amplitudes $a_{j\sigma n}$ and $b_{j\sigma n}$ are defined in the eigenvalue equation $H\phi_{n}^{(r)}=\epsilon_{n}\phi_{n}^{(r)}$ with the $n$th right column eigenvector $\phi_{n}^{(r)}=(\{a_{j\uparrow n},a_{j\downarrow n},b_{j\uparrow n},b_{j\downarrow n}\})^{T}$. Figure 3(c) shows the averaged density distributions $(\bar{\rho}_{j\uparrow}/\bar{\rho}_{1\uparrow},\bar{\rho}_{j\downarrow}/\bar{\rho}_{L\downarrow})$ normalized by the corresponding localized boundary values versus the scaled unit cell index $j/L$ for different system sizes. The collapse into the same lines for respectively upper and lower legs verify the scale-free power law with the power $c=1$.
In addition, the critical NHSEs also have impact on the end states. Since the zero-mode end state protected by the chiral symmetry appears under OBCs when $\delta^{2}<$ ($>$) $\kappa^{2}-1$ for the nonreciprocal SSH chain with (non-)reciprocal unit cells being the boundary cells [18], in the decoupled case the two relatively shifted nonreciprocal SSH chains as a whole always have end states under OBCs with an arbitrary kind of boundary cells. This means that the topological end states in small-sized ladder with weak inter-leg couplings under OBCs may evolve into the bulk states of a topologically trivial phase in the thermodynamic limit, as shown in the insets of Fig. 3(b). Now one may notice that the topological phase diagram in Fig. 1(b) is singular at $\chi^{2}=0$ (the gray vertical line) in the thermodynamic limit.
It is apparent that if we add a leg-dependent potential, the NHSE under OBCs along with the breakdown of BBC will return in the thermodynamic limit, with NHSEs depending on the eigenenergies, and thus the so-called bipolar NHSE will be naturally realized [58, 60].
VI Conclusion
In conclusion, we thoroughly investigate the non-Abelian geometrical, topological and dynamical properties of an effective non-Hermitian description of a dissipative two-leg SSH ladder, i.e., a ladder of coupling two relatively shifted SSH chains with opposite nonreciprocal hoppings. We give the exact topological phase diagram characterized by defining a gauge-invariant winding number under PBCs for chiral symmetric multi-band systems, along with the various geometries of band structure based on the underlying symmetries of the ladder. The evolution of a Bloch state under an external constant force is also investigated in the pseudo-Hermitian symmetric regime to reveal the non-Abelian dynamics in the non-Hermitian multi-band subspace. Under OBCs, the BBC is found conserved for the non-Abelian topology in the thermodynamic limit of the ladder, but the critical NHSEs lead to the leg-dependent boundary localization of bulk states in finite-sized systems, obeying the scale-free law of the decaying length $\xi(L)\propto L$.
Acknowledgements.
This work was supported by the National Key Research and Development Program of China (Grant No. 2022YFA1405304), the Guangdong Basic and Applied Basic Research Foundation (Grant No. 2024A1515010188), the National Natural Science Foundation of China (Grant No. 11904109), and the Startup Fund of South China Normal University.
Appendix A Recap of the pseudo-Hermitian symmetry
To be self-consistent, we recap the concept of pseudo-Hermiticition symmetry in this section. Here, we assume that $\hat{H}$ is a diagonalizable linear operator acting on the Hilbert space of dimension $N$, which has a set of complete biorthonormal right/left eigenvectors $\{|\phi^{(r/l)}_{n}\rangle\}$ and a discrete spectrum $\{\epsilon_{n}\}$:
$\hat{H}$ $=$ $\sum_{n=1}^{N}\epsilon_{n}|\phi^{(r)}_{n}\rangle\langle\phi^{(l)}_{n}|.$ (35)
@@ Appendix A Recap of the pseudo-Hermitian symmetry
00 0 $=$ $\langle\phi_{n}^{(r)}|\hat{H}^{\dagger}\hat{\eta}|\phi_{n}^{(r)}\rangle-\langle\phi^{(r)}|\hat{H}^{\dagger}\hat{\eta}|\phi_{n}^{(r)}\rangle$ (38)
@@
$\hat{H}$ $=$ $\sum_{\text{Im}\,\epsilon_{n}=0}\epsilon_{n}|\phi^{(r)}_{n}\rangle\langle\phi^{(l)}_{n}|$ (39)
$+\sum_{\text{Im}\,\epsilon_{n}>0}\big{(}\epsilon_{n}|\phi^{(r)}_{n}\rangle\langle\phi^{(l)}_{n}|+\epsilon_{n}^{*}|\tilde{\phi}^{(r)}_{n}\rangle\langle\tilde{\phi}^{(l)}_{n}|\big{)},$
@@
$\hat{\eta}$$\equiv$ $\sum_{\text{Im}\,\epsilon_{n}=0}\sigma_{n}|\phi^{(l)}_{n}\rangle\langle\phi^{(l)}_{n}|$
$+\sum_{\text{Im}\,\epsilon_{n}>0}\big{(}|\tilde{\phi}^{(l)}_{n}\rangle\langle\phi^{(l)}_{n}|+|\phi^{(l)}_{n}\rangle\langle\tilde{\phi}^{(l)}_{n}|\big{)},$
$\hat{\eta}^{-1}$$\equiv$ $\sum_{\text{Im}\,\epsilon_{n}=0}\sigma_{n}|\phi^{(r)}_{n}\rangle\langle\phi^{(r)}_{n}|$ (40)
$+\sum_{\text{Im}\,\epsilon_{n}>0}\big{(}|\phi^{(r)}_{n}\rangle\langle\tilde{\phi}^{(r)}_{n}|+|\tilde{\phi}^{(r)}_{n}\rangle\langle\phi^{(r)}_{n}|\big{)},$
@@
which is just the $O^{\dagger}O$-pseudo-Hermiticity of $\hat{H}$. This means that the operator with real spectrum must be $O^{\dagger}O$-pseudo-Hermitian. The inverse statement can also be easily proved. Thus, the necessary and sufficient condition for an operator with real spectrum is the $O^{\dagger}O$-pseudo-Hermiticity with $\hat{O}$ being an inversible linear operator, that is, there must be a spectial $\eta$-pseudo-Hermiticity of $\hat{H}$ with $\hat{\eta}=\hat{O}^{\dagger}\hat{O}$ [13].
@@
with $\xi_{n}>0$ being the eigenvalues and $|\xi_{n}\rangle$ the orthonormal eigenvectors. The quasi-Hermiticity Eq. (45) implies that $\hat{H}$ has the real spectrum, that is, it is similar to a Hermitian operator $\hat{H}_{h}=\hat{S}\hat{H}\hat{S}^{-1}$ that has the identical spectrum, where the similarity operator $\hat{S}$ can be constructed as
@@
where we use the quasi-Hermiticity of $\hat{H}$ in Eq. (45). Thus, quasi-Hermiticity is a sufficient yet necessary condition for the reality of the spectrum.
Appendix B The symmetries of the effective non-Hermitian Hamiltonian
In this section, we give the explicit forms of various symmetries of the effective non-Hermitian Hamiltonian $\hat{H}_{\text{NH}}$ in the main text when $\chi_{+}=\chi_{-}\equiv\chi$:
@@ Appendix B The symmetries of the effective non-Hermitian Hamiltonian
where the bases in respective real and $k$-spaces read
@@
Due to the reality of $\hat{H}_{\text{HN}}$’s parameters, the ladder respects the time-reversal symmetry:
$\hat{T}\hat{H}_{\rm NH}\hat{T}^{-1}=\hat{H}_{\rm NH}$ (52)
where the time-reversal operation $\hat{T}$ acts as the complex conjugation $K$ for numbers, preserving the operators in real space:
$\hat{T}:\hat{a}_{j\sigma}^{({\dagger})}\rightarrow\hat{a}_{j\sigma}^{({\dagger})}K,\,\hat{b}_{j\sigma}^{({\dagger})}\rightarrow\hat{b}_{j\sigma}^{({\dagger})}K,$ (53)
and correspondingly in $k$-space, it reverses the sign of $k$ accompanied by the complex conjugation:
$\hat{T}:\hat{a}_{k\sigma}^{({\dagger})}\rightarrow\hat{a}_{-k,\sigma}^{({\dagger})}K,\,\hat{b}_{k\sigma}^{({\dagger})}\rightarrow\hat{b}_{-k,\sigma}^{({\dagger})}K.$ (54)
@@
$KH_{k}K=H_{k}^{*}=H_{-k},$ (56)
which ensures that the eigenenergies in $k$-space have the relation $\epsilon_{-k,m}=\epsilon_{kn}^{*}$, where $\epsilon_{kn}$ means the $n$-th eigenenergy of $H_{k}$.
@@
which indicates that $\hat{C}^{\dagger}=\hat{C}^{-1}=\hat{C}$. Hereafter, $\tau_{x,y,z}$ and $\sigma_{x,y,z}$ are both Pauli matrices acting on the sublattice ($a,b$) and the leg ($\uparrow$,$\downarrow$ , ↓) spaces, respectively; $\tau_{0}$ and $\sigma_{0}$ are the respective identity matrices. The chiral symmetry Eq. (57) requires that
@@
which ensures that the eigenenergies in $k$-space have the relation $\epsilon_{km}=-\epsilon_{kn}$.
@@
which ensures that the eigenenergies in $k$-space have the relation $\epsilon_{-k,m}=\epsilon_{kn}^{*}$.
Inversion symmetry.
The ladder also has the inversion symmetry:
$\hat{I}\hat{H}_{\rm NH}\hat{I}^{-1}=\hat{H}_{\rm NH}$ (69)
with the inversion operation
in real space, where $\bar{\sigma}$ means the opposite leg to leg-$\sigma$, and correspondingly
in $k$-space. Thus,
with
$I_{k}^{\dagger}$ $=$ $I_{k}^{-1}\neq I_{k},$ (80)
$I_{k}H_{k}I_{k}^{\dagger}=H_{-k},$ (81)
which ensures that the eigenenergies in $k$-space have the relation $\epsilon_{-k,m}=\epsilon_{kn}$.
@@
which ensures that the eigenenergies in $k$-space have the relation $\epsilon_{km}=\epsilon_{kn}^{*}$.
@@
in $k$-space, which indicates that $\hat{\eta}^{\dagger}=\hat{\eta}^{-1}=\hat{\eta}$. Correspondingly, using the discrete Fourier transform (6), we obtain the operations in real space:
@@
When $\chi_{+}\neq\chi_{-}$, one can follow the similar procedure to verify that the time-reversal, the chiral, and the inversion symmetries are preserved, but the pseudo-mirror, the pseudo-glide, and the $\eta_{k}$-pseudo-Hermitian symmetries are broken. Even in this case, the preserving symmetries are enough to ensure the $D_{2h}$ point group of the $H_{k}$’s eigenenergies in the composite 3D space spanned by the three 2-fold rotating axies of the $k$-axis, the real and the imaginary axes of the complex energy plane.
Appendix C A gauge-invariant winding number for non-Hermitian multi-band systems with chiral symmetries
Due to the chiral symmetry $\{H_{k},C\}=0$, the right/left eigenvectors of the Hamiltonian matrix $H_{k}$ of dimension $2N$ appear in pairs for each $k$, $\{Cu^{(r/l)}_{kn},u^{(r/l)}_{kn}\}$, with the corresponding sign-inverted eigenenergies $\{\epsilon_{kn},-\epsilon_{kn}\}$. Thus, we can categorize these eigenvectors into two subspaces $\mathbb{G}_{\pm}$ of dimension $N$, and define a $2N\times N$ matrix $U_{-}^{(r/l)}\equiv(u_{k1}^{(r/l)},\cdots,u_{kN}^{(r/l)})$ to group the eigenvectors in $\mathbb{G}_{-}$ and the matrix $U_{+}^{(r/l)}\equiv CU_{-}^{(r/l)}$ in $\mathbb{G}_{+}$, where the subscripts “$\pm$” are simply labels, as the eigenenergies are generally complex for non-Hermitian Hamiltonians. One can choose eigenvectors of the subspaces for demand, without necessarily ordering them by the real or imaginary parts of the eigenenergies. Hereafter, we assume that the defectiveness (i.e., non-diagonalization) of the non-Hermitian Hamiltonian $H_{k}$ occurs only at finite discrete sets of system parameters such that the biorthogonal theory can be analytically used at the neighborhood of these points, preventing using the theory of generalized eigenvectors.

### v2

@@ V Non-Abelian dynamics in the pseudo-Hermitian symmetric regime
$\alpha(t)$ $\approx$ $\mathcal{P}\exp\Big{[}-i\int_{k_{0}}^{k}dkA_{k}\Big{]}\alpha(0)$ (31)
$\equiv$ $W_{k_{0}\rightarrow k}\alpha(0)=U_{k}^{-1}U_{k_{0}}\alpha(0),$
where
$W_{k_{0}\rightarrow k}\equiv\mathcal{P}\exp\Big{[}-i\int_{k_{0}}^{k}dkA_{k}\Big{]}$ (32)
is a four-band Wilson line for non-Hermitian systems along the path from $k_{0}$ to $k$ in $k$ space [50]. Equation (31) means that the evolution only depends on the start and the final points, regardless of the path’s selection, as shown in Fig. 3(h), which reproduces the evolution in Fig. 3(g) by Eq. (29) under a very strong force. Note that after one BZ cycle, the state will totally return to the initial state, which can be directly proved by noting that four-band Wilson line in Eq. (31) becomes identity.
When $w_{2}\ll F\ll\Delta$, where $w_{2}$ is the maximum energy difference between the lowest (highest) two bands and $\Delta$ is the minimum energy difference (i.e., the gap) between the middle two bands, the lowest (highest) two bands can be approximated as being degenerate in the whole BZ and the excitation to the other two bands can be ignored, that is, the dynamics can be considered as an adiabatic evolution in the subspace of the lowest (highest) two degenerate bands. Thus, Eq. (29) can be approximately decoupled into two independent evolutions in each subspace:
$\alpha^{(\pm)}(t)$ $\approx$ $\mathcal{P}\exp\Big{[}-i\int_{k_{0}}^{k}dkA^{(\pm)}_{k}\Big{]}\alpha^{(\pm)}(0)$ (33)
$\equiv$ $W^{(\pm)}_{k_{0}\rightarrow k}\alpha^{(\pm)}(0),$
where the +()+~(-) labels the upper (lower) $2\times 2$ diagonal block for the labeled matrices or the upper (lower) $2\times 1$ block for the labeled column vectors. In this case, the non-Abelian dynamics in corresponding subspaces can be well described by two-band Wilson lines for non-Hermitian systems [50]:
$W^{(\pm)}_{k_{0}\rightarrow k}\equiv\mathcal{P}\exp\Big{[}-i\int_{k_{0}}^{k}dkA^{(\pm)}_{k}\Big{]}$, (34)
as shown in Fig. 3(i), which roughly captures the qualitative feature of the direct dynamics in Fig. 3(c) calculated by Eq. (29). The quantitative difference between Figs. 3(c) and 3(i) mainly comes from the dissatisfaction of the lower-bound condition $F\gg w_{2}$. Different from the four-band Wilson line, the state evolved by the two-band Wilson line in Eq. (33) does not return to its initial state after one BZ cycle due to the crossing point in the spectrum. The details of the numerical method for calculating the two Wilson lines can be referred to in Appendix D.
The non-Abelian dynamics for other strengths of the force in Fig. 3 cannot be well captured by the two Wilson lines. However, the common feature is that the population in each subspace smoothly transfers from one band to the other across the degenerate point, manifested by the sudden change of the auxiliary lines in Figs. 3(b)–3(i). The similar phenomenon also occurs in the non-Abelian dynamics of Hermitian multi-band systems [9].
VI Conclusion and discussion
In conclusion, we investigate the non-Abelian properties of the nonreciprocal SSH ladder that is formed by coupling two nonreciprocal SSH chains. Under PBCs, we exactly obtain the phase diagram of the geometry of band structure, including the real and complex bands, as well as the real line, imaginary line, and point gaps, and also obtain the phase diagram of the non-Abelian topology under chiral symmetry, characterized by a newly defined gauge-invariant winding number, which is different from the common-used ones. Under OBCs, the BBC is found survived in the thermodynamic limit, but breaks down for finite sizes when the inter-leg coupling is weak, where the critical NHSE emerges, obeying the scale-free power law $\xi(L)\propto L$. The non-Abelian dynamics of a Bloch state subject to an external constant force in the pseudo-Hermitian symmetric regime, where all bands are real, is also studied in comparison with the two-band and four-band Wilson lines.
In experiments, this nonreciprocal SSH ladder can be potentially realized using a time-multiplexed photonic resonator network as in Ref. [49] instead by coupling two SSH chains and replacing the jump operators therein by Eq. (3). Alternatively, one may also simulate this ladder using ultracold atoms with the technique of nonreciprocal transport in momentum space [62, 63]. The classical electric circuit is also a good candidate for this ladder [60, 64].
Acknowledgements.
We thank L. He for the helpful discussion on the explicit expression of the psudo-Hermitian operator in real space. This work was supported by the National Key Research and Development Program of China (Grant No. 2022YFA1405304), the Guangdong Basic and Applied Basic Research Foundation (Grant No. 2024A1515010188), the National Natural Science Foundation of China (Grant No. 11904109), and the Startup Fund of South China Normal University.
data availability
The data that support the findings of this article are not publicly available. The data are available from the authors upon reasonable request.
Appendix A Recap of the pseudo-Hermitian symmetry
To be self-consistent, we recap the concept of pseudo-Hermitian symmetry in this section. Here, we assume that $\hat{H}$ is a diagonalizable linear operator acting on the Hilbert space of dimension $N$, which has a set of complete biorthonormal right and left eigenvectors $\{|\phi^{(r/l)}_{n}\rangle\}$ and a discrete spectrum $\{\epsilon_{n}\}$:
@@ Appendix A Recap of the pseudo-Hermitian symmetry
$0$ $=$ $\langle\phi_{n}^{(r)}|\hat{H}^{\dagger}\hat{\eta}|\phi_{n}^{(r)}\rangle-\langle\phi^{(r)}|\hat{H}^{\dagger}\hat{\eta}|\phi_{n}^{(r)}\rangle$ (38)
@@
[…]
@@
[…]
@@
which is just the $O^{\dagger}O$-pseudo-Hermiticity of $\hat{H}$. This means that the operator with real spectrum must be $O^{\dagger}O$-pseudo-Hermitian. The inverse statement can also be easily proved. Thus, the necessary and sufficient condition for an operator with real spectrum is the $O^{\dagger}O$-pseudo-Hermiticity with $\hat{O}$ being an invertible linear operator, that is, there must be a special $\eta$-pseudo-Hermiticity of $\hat{H}$ with $\hat{\eta}=\hat{O}^{\dagger}\hat{O}$ [13].
@@
with $\xi_{n}>0$ being the eigenvalue of the orthonormal eigenvector $|\xi_{n}\rangle$. The quasi-Hermiticity is a sufficient yet necessary condition for the reality of the spectrum [13]. For the invertible case of $\hat{\xi}$, it means that there exists a Hermitian operator $\hat{H}_{h}=\hat{S}\hat{H}\hat{S}^{-1}$ that has the identical spectrum, where the similarity operator $\hat{S}$ can be constructed as
@@
where we use the quasi-Hermiticity of $\hat{H}$ in Eq. (45).
Appendix B Symmetries of $\hat{H}_{\text{NH}}$
In this section, we give the explicit forms of symmetries for the non-Hermitian Hamiltonian $\hat{H}_{\text{NH}}$ in the main text when $\chi_{+}=\chi_{-}\equiv\chi$:
@@ Appendix B Symmetries of $\hat{H}_{\text{NH}}$
where the bases in respective real and $k$ spaces read as
@@
Due to the reality of all parameters in $\hat{H}_{\text{NH}}$, the ladder respects the time-reversal symmetry:
$\hat{T}\hat{H}_{\rm NH}\hat{T}^{-1}=\hat{H}_{\rm NH},$ (52)
where the time-reversal operator $\hat{T}$ acts just as the complex conjugation on numbers, preserving the operators in real space,
$\hat{T}:\hat{a}_{j\sigma}^{({\dagger})}\rightarrow\hat{a}_{j\sigma}^{({\dagger})},~~\hat{b}_{j\sigma}^{({\dagger})}\rightarrow\hat{b}_{j\sigma}^{({\dagger})},~~i\rightarrow-i,$ (53)
and correspondingly in $k$ space, it reverses the sign of $k$ of the operators:
$\hat{T}:\hat{a}_{k\sigma}^{({\dagger})}\rightarrow\hat{a}_{-k,\sigma}^{({\dagger})},~~\hat{b}_{k\sigma}^{({\dagger})}\rightarrow\hat{b}_{-k,\sigma}^{({\dagger})}.$ (54)
@@
$H_{k}^{*}=H_{-k},$ (56)
which ensures that the eigenenergies in $k$ space have the relation $\epsilon_{-k,m}=\epsilon_{kn}^{*}$, where $\epsilon_{kn}$ means the $n$-th eigenenergy of $H_{k}$.
@@
which indicates that $\hat{C}^{\dagger}=\hat{C}^{-1}=\hat{C}$. Hereafter, $\tau_{x,y,z}$ and $\sigma_{x,y,z}$ are both Pauli matrices acting on the sublattice ($a,b$) and the leg ($\uparrow,\downarrow$) spaces, respectively; $\tau_{0}$ and $\sigma_{0}$ are the respective identity matrices. The chiral symmetry Eq. (57) requires that
@@
which ensures that the eigenenergies in $k$ space have the relation $\epsilon_{km}=-\epsilon_{kn}$.
Inversion symmetry.
The ladder also has the inversion symmetry:
$\hat{I}\hat{H}_{\rm NH}\hat{I}^{-1}=\hat{H}_{\rm NH}$ (63)
with the inversion operation
in real space, where $\bar{\sigma}$ means the opposite of leg $\sigma$, and correspondingly
in $k$ space. Thus,
with
$I_{k}^{\dagger}$ $=$ $I_{k}^{-1}\neq I_{k},$ (74)
$I_{k}H_{k}I_{k}^{\dagger}=H_{-k},$ (75)
which ensures that the eigenenergies in $k$ space have the relation $\epsilon_{-k,m}=\epsilon_{kn}$.
@@
which ensures that the eigenenergies in $k$ space have the relation $\epsilon_{-k,m}=\epsilon_{kn}^{*}$. This symmetry can be regarded as a non-Hermitian generalization of the traditional mirror symmetry in Hermitian systems.
@@
which ensures that the eigenenergies in $k$ space have the relation $\epsilon_{km}=\epsilon_{kn}^{*}$. This symmetry can be regarded as a non-Hermitian generalization of the traditional glide symmetry in Hermitian systems.
@@
in $k$ space, which indicates that $\hat{\eta}^{\dagger}=\hat{\eta}^{-1}=\hat{\eta}$. Correspondingly, using the discrete Fourier transform (7), we obtain the operations in real space:
@@
Compared with the Hamiltonian $\hat{H}$ of the Hermitian SSH ladder, the dissipation-induced nonreciprocal hopping in $\hat{H}_{\text{NH}}$ relaxes some traditional symmetries, such as the mirror and the glide symmetries, to the corresponding pseudo ones, and also generalizes the time-reversal symmetry to the field of complex energies.
When $\chi_{+}\neq\chi_{-}$, one can follow the similar procedure to verify that the time-reversal, the chiral, and the inversion symmetries are preserved, but the pseudo-mirror, the pseudo-glide, and thus the $\eta_{k}$-pseudo-Hermitian symmetries are broken. Even in this case, the preserved symmetries are sufficient to ensure the $D_{2h}$ symmetry of the band structure in the composite 3D space spanned by the three twofold rotating axies: the $k$ axis, the real and the imaginary axes of the complex energy plane. This symmetry implies that the ladder possesses other hidden pseudo-Hermitian symmetry instead of the $\eta_{k}$-pseudo-Hermitian symmetries in Eq. (9) according to the statement in Appendix A, although the explicit form of the pseudo-Hermitian operator cannot be easily written.
Appendix C A gauge-invariant winding number for non-Hermitian multi-band systems under the chiral symmetry
Due to the chiral symmetry $\{H_{k},C\}=0$, the right or left eigenvectors of the Hamiltonian matrix $H_{k}$ of dimension $2N$ appear in pairs for each $k$, $\{Cu^{(r/l)}_{kn},u^{(r/l)}_{kn}\}$, with the corresponding sign-inverted eigenenergies $\{\epsilon_{kn},-\epsilon_{kn}\}$. Thus, we can categorize these eigenvectors into two subspaces $\mathbb{G}_{\pm}$ of dimension $N$, and define a $2N\times N$ matrix $U_{-}^{(r/l)}\equiv(u_{k1}^{(r/l)},\dots,u_{kN}^{(r/l)})$ to group the eigenvectors in $\mathbb{G}_{-}$ and the matrix $U_{+}^{(r/l)}\equiv CU_{-}^{(r/l)}$ in $\mathbb{G}_{+}$, where the subscripts $\pm$” are simply labels, as the eigenenergies are generally complex for non-Hermitian Hamiltonians. One can choose eigenvectors of the subspaces for demand, without necessarily ordering them by the real or imaginary parts of the eigenenergies. Hereafter, we assume that the defectiveness (i.e., non-diagonalization) of the non-Hermitian Hamiltonian $H_{k}$ occurs only at finite discrete sets of system parameters such that the biorthogonal theory can be analytically used at the neighborhood of these points, preventing using the theory of generalized eigenvectors.